# Sea ice floe segmentation in close-range optical imagery using active contour and foundation models

### v3

@@ Abstract
The size of sea ice floes is a key factor influencing ice coverage, albedo, wave propagation through ice-covered waters, and ocean-atmosphere energy exchanges. Floe size can be observed by processing visual-range imagery from ships, aircraft, or satellites. However, autonomously capturing floe boundaries from imagery remains challenging, particularly due to the heterogeneity of sea ice, which impairs boundary definition and reduces image clarity. This study evaluates the accuracy of sea ice floe segmentation using the gradient vector flow (GVF) algorithm, the deep learning-based Segment Anything Model (SAM), and a hybrid approach combining GVF and SAM. These methods are applied to the most extensive dataset of close-range, high-resolution sea ice imagery currently available, collected from cameras aboard an icebreaker during an Antarctic winter expedition. Spanning a wide range of ice conditions and image clarity, the dataset provides a rigorous test of segmentation approaches. Their performance is assessed in terms of floe detection accuracy, floe size distribution, and ice concentration, with results compared against a manually segmented benchmark. The findings indicate that SAM, when used in prompt-driven mode, offers the best balance between accuracy and computational efficiency. Its strong performance in estimating sea ice concentration and detecting floes, while maintaining close agreement with benchmark floe size distributions, makes it suitable for real-time applications and scalable analysis of large imagery datasets.
@@ 1 Introduction
Observations of floe geometry, coverage, size, and dynamics are key to advancing and validating sea ice models (e.g., Buckley et al., 2024; Roach et al., 2018). Their characterization can be inferred from high-resolution imagery acquired via satellite, airborne, and shipborne sensors, combined with image processing techniques (e.g., Alberello et al., 2022; Alsharay et al., 2022; Herman et al., 2021; Lu et al., 2016; Paget et al., 2001; Parmiggiani et al., 2018; Sandru et al., 2020; Toyota et al., 2006; Zhang and Hughes, 2023; Zhang and Skjetne, 2015; Weissling et al., 2009). Traditional methods, such as threshold segmentation, rely on user-defined values, which can introduce inconsistencies due to variations in sea ice conditions across the image (Heyn et al., 2017; Lu et al., 2008; Alberello et al., 2019). To address this limitation, more advanced techniques, including k-means clustering (Weissling et al., 2009; Lu et al., 2016), the gradient vector flow (GVF) snake algorithm (Zhang and Skjetne, 2015), and watershed segmentation (Dumas-Lefebvre and Dumont, 2023; Parmiggiani et al., 2018), have been applied to improve automation and robustness in sea ice analysis. Despite their advantages, these methods face challenges when applied to complex atmospheric and ice conditions that reduce the clarity of sea ice imagery. Although they perform well under optimal conditions, accuracy and consistency decline markedly in images with reduced clarity. Factors such as coarse resolution, cloud cover, fog, variable lighting, and low contrast between ice and water commonly reduce image clarity. Additionally, snow cover and the presence of interstitial ice, thin ice forming in the spaces between floes, can obscure floe boundaries, creating the illusion of a more uniform surface or larger floes. These challenges complicate the identification of individual floes, particularly when they are densely packed, and make automated analysis of large image datasets particularly problematic.
@@
Here, we use an image acquisition protocol deployed on the icebreaker S.A. Agulhas II during expeditions to the Antarctic marginal ice zone (MIZ) in the austral winter (Alberello et al., 2019, 2022; Toffoli et al., 2024). The resulting dataset, the most extensive currently available, comprises high-resolution, close-range images of sea ice that offer detailed spatial and temporal coverage, capturing a broad range of ice conditions under varying levels of clarity. Taking advantage of this diversity, we assess the robustness of different segmentation approaches, namely the traditional active contours model (GVF Snake algorithm), the advanced SAM deep learning model, and a novel integration of the two. To support this evaluation, we develop a fully automated processing pipeline that enables their systematic and autonomous application across the varied ice conditions represented in the dataset. We conduct detailed, floe-by-floe visual comparisons against manually segmented benchmarks for two representative scenes with contrasting ice conditions, revealing the segmentation discrepancies, strengths, and limitations of each method. Beyond qualitative assessment, we perform a statistical evaluation using a benchmark of over 2,800 manually segmented floes, the most extensive dataset of its kind for sea ice segmentation analysis, to identify the most effective method for detecting ice floes and estimating floe size and ice concentration under diverse conditions.
@@ 2 Image acquisition
Figure 1: Antarctic SCALE expedition, winter 2022: (a) camera mounted on S.A. Agulhas II; (b) ship track from departure on 11 July 2022 to return on 31 July 2022, entering the sea ice region on 19 July and exiting on 24 July; (c–e) high-resolution images of the sea ice surface captured under (c) conditions featuring distinct ice floes with defined edges, (d) diffuse sea ice boundaries characterized by densely packed, irregularly shaped floes with indistinct or visually merged edges, and (e) at nighttime.
@@
The sensor consisted of a GigE monochrome industrial camera (Fig. 1a) equipped with 2/3” Sony CMOS Pregius sensor and 5 mm F1.8 C-mount lenses (angle of view $\approx$120). It was mounted on the port side of the monkey bridge at 25 m above the water line and tilted by $\approx$70 relative to the ground. The field of view covered a portion of the ocean surface of $\approx$ 200$\times$200 m2. Images were recorded continuously with a resolution of 1920$\times$1080 pixels and at a sampling rate of 1 Hz. The sensor was paired with a synchronized inertial measurement unit (IMU) to record variations in the camera’s position relative to the horizontal plane during navigation.
Around 90,000 images were collected during the expedition, covering both day and night, with nighttime imaging supported by flashlights illuminating the sea ice (see sample images in Fig. 1c–e). Consequently, the database spans a wide range of ice and atmospheric conditions, with approximately 70% capturing complex sea ice scenarios. These include low light conditions (Fig. 1e) and densely packed, irregularly shaped floes with diffuse boundaries, where floe edges are indistinct or merge visually due to snow cover, interstitial ice, or low contrast (Fig. 1d). This diversity offers a robust testbed for evaluating image processing methods.
@@ 3.1 Orthorectification
Orthorectification interpolates over the pixels in the original image based on the camera projection matrix, which maps the conversion of a three-dimensional point cloud in the real world to a two-dimensional plane (the image) through intrinsic and extrinsic parameters (Hartley and Zisserman, 2004). The former depends on how the sensor captures images and include, but are not limited to, the focal length, aperture, resolution, and the optical center of the camera’s sensor. These parameters are derived from a calibration process (e.g., Zhang, 2000), which evaluates image distortion by analyzing multiple images of a planar pattern with known geometry (e.g., a chessboard; see Sturm and Maybank, 1999; Bouguet, 2004). Calibration was performed before and after the expedition to verify that intrinsic features remain constant throughout the campaign.
@@
Floe edges lose sharpness with distance from the sensor (see upper band of Fig. 2b). To avoid ambiguity in the far field and, hence, reduce the risk of detecting unrealistic floes, the workable portion of the orthorectified images is restricted to an area of about 60 m $\times$ 120 m close to the sensors (see area within the red rectangle in Fig. 2b).
@@ 3.2 Effect of ship motion
The assumption that the sensor remains fixed relative to the ocean surface holds true in calm seas but does not apply in open ocean conditions (cf. Alberello et al., 2022). Movement of the supporting platform can alter the camera’s orientation affecting orthorectification and floe size dimensions. To assess sensitivity to ship motion, an image captured in clear sky conditions was used as a proxy. The orthorectification was evaluated by artificially introducing misalignments of $\pm$1 m in elevation, $\pm$1 in tilt, and $\pm$2 in roll relative to the reference configuration established during calm seas. These values were chosen as representative of extreme ship motions observed during the expedition. A few clearly distinguishable floes were manually segmented after orthorectification by outlining their contours using photo editing software for analysis (see Fig. 3).
@@ 3.3 Image enhancement
For night photographs, enhancement was limited to areas illuminated by the vessel’s spotlight. These regions are identified using haze removal techniques (He et al., 2011), which estimate the illumination map by inverting the low-light image, applying dehazing, and reinverting the result. This highlights the spotlight-illuminated areas for further analysis.
@@ 4 Image segmentation
4.1 Active contours model
@@
The adjacent objects are then separated by smoothing their edges through a morphological erosion process (Adams, 1993). The newly formed objects are initialized using ellipses, which approximate the shape of ice floes (Alberello et al., 2019; Passerotti et al., 2022). The initial contour is then chosen such that the second moment of the covariance matrix (a 2$\times$2 matrix that captures the object’s spatial spread and orientation) aligns with that of the object (Mulchrone and Choudhury, 2004) (see Fig. 4d).
@@
The GVF snake algorithm produces anomalies and inaccuracies, which can result in features that do not exist in the original image (“hallucinations”). To quantify how closely the contours resemble a perfect circle or how elongated the features are, coefficients of circularity and eccentricity are used (e.g., Ayoub, 2003; Bottema, 2000; Alberello et al., 2019). By applying high-valued thresholds, spurious features in the form of nearly perfect circles or line segments are removed (circularity >1 and eccentricity $\geq$0.9). Additionally, objects with open boundaries at the image’s edges are excluded from the segmentation.
@@
From the final binary mask, each ice floe is identified as a connected region of pixels using eight-neighborhood connectivity, meaning that two pixels are considered connected if they share an edge or a corner. Once the floes are identified, geometric properties such as area, perimeter, and diameter are derived. The area is determined by counting the total number of pixels in each region. The perimeter is obtained by tracing along the region boundary to measure its continuous outline (Biswas and Hazra, 2018). The diameter is calculated by equating the floe to a circle of the same area, thus representing the diameter of that hypothetical circle. The ice concentration is computed by summing the areas of all floes and dividing by the total image size (excluding any masked-out, empty regions introduced by orthorectification).
@@ 4.2.1 Automatic mask generator (SAM-A)
Figure 5: SAM’s automatic segmentation mode (SAM-A): (a) grid-based fixed points automatically selected by SAM on the enhanced sea ice image; (b) consolidation process applied to the list of masks returned by the model after filtering; (c) final unified binary mask containing all detected floes.
@@
SAM cannot differentiate between the various segmented elements and, hence, segmentation includes both ice floes and open water fractions. The latter is excluded by applying the convexity coefficient (Sklansky, 1970) as a filtering criterion. The complexity coefficient quantifies how closely an object’s boundary resembles that of a convex shape. Ice floes, with regular boundaries that can often be approximated by ellipses or circles (Alberello et al., 2019), typically exhibit convexity values close to 1. In contrast, open water regions, with more irregular boundaries, have lower convexity values. A cut-off value of 0.8 is set, such that masks with a convexity less than or equal to 0.8 are excluded (cf. Wang et al., 2024). The remaining masks (floes) are analyzed to identify those with anomalous shapes or open boundaries at the edges of the image using eccentricity and circularity metrics (see §4.1). Masks exhibiting these irregularities are also eliminated because they are either inconsistent with typical floe geometry or incompletely segmented due to edge effects.
@@ 4.2.2 Prompt-based (SAM-P)
Figure 6: SAM’s prompt-based segmentation mode (SAM-P): (a) seed points (red dots) marking floe locations on the enhanced sea ice image; (b) consolidation process applied to the list of masks returned by the model after filtering; (c) final unified binary mask containing all detected floes.
In prompt-based mode, segmentation is guided by user-defined seeds, which are-derived from the distance transform matrix introduced in §4.1 (Fig.4b). These seeds pre-identify the floes, and, for each one, SAM-P generates a corresponding mask (Fig.6a). In regions with poorly defined boundaries, the assignment of seeds to floes can be ambiguous, and the resulting mask does not always isolate the intended floe accurately. In such cases, the mask includes unwanted areas, such as discontinuous water regions or parts of adjacent floes (see the first two layers in Fig.6b). Shape-based filters for circularity, eccentricity, and convexity, with user-defined thresholds (see §4.1 and §4.2.1), are applied to remove clusters of feature pixels that do not match the typical shape of ice floes. A union operation is also used to merge clusters representing parts of adjacent floes with the correct floe. This operation combined all the masks in the output list into a single unified mask, automatically consolidating the segmented portions of floes into their appropriate masks (Fig.6c). From this final binary mask, the properties of each individual ice floe and sea ice concentration are extracted.

### v4

@@ Abstract
The size of sea ice floes in the marginal ice zone (MIZ) is a key factor influencing ice coverage, albedo, wave propagation through ice-covered waters, and ocean–atmosphere energy exchanges. Floe size can be observed by processing visual-range imagery from ships, aircraft, or satellites. However, autonomously capturing floe boundaries from imagery remains challenging, particularly due to the heterogeneity of sea ice, which impairs boundary definition and reduces image clarity. This study evaluates the accuracy of sea ice floe segmentation using the gradient vector flow (GVF) active contour method, the deep learning-based Segment Anything Model (SAM), and a hybrid approach combining GVF and SAM. These methods are evaluated on a representative subset of a large dataset of close-range, high-resolution sea ice imagery, collected from cameras aboard an icebreaker during an Antarctic winter expedition. Spanning a wide range of ice conditions and image clarity in the MIZ, the subset provides a rigorous test bed of segmentation approaches. Their performance is assessed in terms of floe detection accuracy, floe size distribution, and ice concentration, with results compared against a manually segmented benchmark. The findings indicate that SAM, when used in prompt-driven mode, offers the best balance between accuracy and computational efficiency. Its strong performance in estimating sea ice concentration and detecting floes, while maintaining close agreement with benchmark floe size distributions, makes it suitable for real-time applications and scalable analysis of large imagery datasets.
@@ 1 Introduction
Accurate observations of floe geometry, coverage, size, and dynamics are key to advancing and validating modern sea ice models (e.g., Bennetts et al., 2017; Roach et al., 2018; Wang et al., 2023; Buckley et al., 2024; Day et al., 2024), as these metrics strongly influence sea ice evolution and related climate feedbacks. Large scale observational datasets capturing spatial and temporal variability across the MIZ are necessary to constrain floe size distribution models, which are sensitive to assumptions about floe geometry and evolution (Bateson et al., 2022; Horvat, 2022). Floe characterization has long relied on high-resolution imagery combined with image processing techniques, and is now routinely extended using satellite, airborne, and shipborne sensors (e.g., Alberello et al., 2022; Alsharay et al., 2022; Herman et al., 2021; Lu et al., 2016; Paget et al., 2001; Parmiggiani et al., 2018; Sandru et al., 2020; Toyota et al., 2006; Zhang and Hughes, 2023; Zhang and Skjetne, 2015; Weissling et al., 2009; Hwang et al., 2017). Manual approaches, such as hand-drawn floe delineation, can be accurate, but they are painstaking and time-intensive, and do not scale to the number of images required for robust statistical analyses (Horvat, 2022). Classical automated methods, such as threshold segmentation, rely on user-defined values, which can introduce inconsistencies due to variations in sea ice conditions across the image (Heyn et al., 2017; Lu et al., 2008; Alberello et al., 2019). To address this limitation, more advanced techniques, including k-means clustering (Weissling et al., 2009; Lu et al., 2016), the gradient vector flow (GVF) snake algorithm (Zhang and Skjetne, 2015), and watershed segmentation (Dumas-Lefebvre and Dumont, 2023; Parmiggiani et al., 2018), have been applied to improve automation and robustness in sea ice analysis. Despite their advantages, these methods face challenges when applied to complex atmospheric and ice conditions that reduce the clarity of sea ice imagery. Although they perform well under optimal conditions, accuracy and consistency decline markedly in images with reduced clarity. Factors such as coarse resolution, cloud cover, fog, variable lighting, and low contrast between ice and water reduce image clarity. Additionally, snow cover and the presence of interstitial ice, thin ice forming in the spaces between floes, can obscure floe boundaries, creating the illusion of a more uniform surface or larger floes. These challenges complicate the identification of individual floes, particularly when they are densely packed, and make automated analysis of large image datasets problematic.
@@
Here, we use an image acquisition protocol deployed on the icebreaker S.A. Agulhas II during expeditions to the Antarctic MIZ in the austral winter (Alberello et al., 2019, 2022; Toffoli et al., 2024). The resulting dataset, the most extensive currently available, comprises high-resolution, close-range images of sea ice that offer detailed spatial and temporal coverage, capturing a broad range of ice conditions under varying levels of clarity. Taking advantage of this diversity, we assess the robustness of different segmentation approaches, namely the traditional active contours model (GVF Snake algorithm), the advanced SAM deep learning model, and a novel integration of the two. To support this evaluation, we develop a fully automated processing pipeline that enables their systematic and autonomous application across the varied ice conditions represented in the dataset. We conduct detailed, floe-by-floe visual comparisons against manually segmented benchmarks for two representative scenes with contrasting ice conditions, revealing the segmentation discrepancies, strengths, and limitations of each method. Beyond qualitative assessment, we perform a statistical evaluation using a benchmark of over 2,800 manually segmented floes, the most extensive dataset of its kind for sea ice segmentation analysis, to identify the most effective method for detecting ice floes and estimating floe size and ice concentration under diverse conditions.
@@ 2 Image acquisition
Figure 1: Antarctic SCALE expedition, winter 2022: (a) camera mounted on S.A. Agulhas II; (b) ship track from departure on 11 July 2022 to return on 31 July 2022, entering the sea ice region on 19 July and exiting on 24 July, overlaid on sea ice concentration from the University of Bremen AMSR2 ASI product (6.25 km grid; Spreen et al., 2008); (c–e) high-resolution images of the sea ice surface captured under (c) conditions featuring distinct ice floes with defined edges, (d) diffuse sea ice boundaries characterized by densely packed, irregularly shaped floes with indistinct or visually merged edges, and (e) at nighttime.
@@
[…]
Around 90,000 images were collected during the expedition, covering both day and night, with nighttime imaging supported by flashlights illuminating the sea ice (see sample images in Fig. 1c–e). Consequently, the database spans a wide range of ice and atmospheric conditions, with approximately 70% capturing complex sea ice scenarios. These situations are characterized by low light (Fig. 1e), densely packed and irregularly shaped floes, and diffuse boundaries, where floe edges are indistinct or visually merge due to snow cover, interstitial ice, or low contrast (Fig. 1d). This diversity offers a robust testbed for evaluating image processing methods
@@ 3.1 Orthorectification
Orthorectification interpolates over the pixels in the original image based on the camera projection matrix, which maps the conversion of a three-dimensional point cloud in the real world to a two-dimensional plane (the image) through intrinsic and extrinsic parameters (Hartley and Zisserman, 2004). The former depends on how the sensor captures images and include, but are not limited to, the focal length, aperture, resolution, and the optical center of the camera’s sensor. These parameters are derived from a calibration process (e.g., Zhang, 2000), which evaluates image distortion by analyzing multiple images of a planar pattern with known geometry (e.g., a chessboard; see Sturm and Maybank, 1999; Bouguet, 2004). Using the method in Zhang (2000), multiple images of a chessboard are captured from various orientations and positions across the field of view. The algorithm automatically detects pattern corners and estimated the intrinsic parameters by fitting the observed geometry to its known dimensions, while correcting for radial and tangential lens distortions. Potential errors, such as imperfect corner localization or slight non-planarity, are minimized by employing a rigid board, good lighting, and a wide range of viewing angles. Parameter drift due to changes in focus or zoom is avoided by keeping lens settings fixed between calibration and deployment. Calibration was performed before and after the expedition to verify that intrinsic features remained stable throughout the campaign.
@@
Floe edges lose sharpness with distance from the sensor (see upper band of Fig. 2b). To avoid ambiguity in the far field and, hence, reduce the risk of detecting unrealistic floes, the workable portion of the orthorectified images is determined by qualitatively assessing multiple images to identify where floe edges noticeably degrade. This effectively restricts the analysis to an area of about 60 m $\times$ 120 m close to the sensor (see area within the red rectangle in Fig. 2b).
@@ 3.2 Effect of ship motion
The assumption that the sensor remains fixed relative to the ocean surface holds true in calm seas but does not apply in open ocean conditions (cf. Alberello et al., 2022). Movement of the supporting platform can alter the camera’s orientation affecting orthorectification and floe size dimensions. To assess sensitivity to ship motion, an image captured in clear sky conditions was used as a proxy. The orthorectification was evaluated by artificially introducing misalignments of $\pm$1 m in elevation, $\pm$1 in tilt, and $\pm$2 in roll relative to the reference configuration established during calm seas. These extreme values of ship motion are determined based on direct measurements obtained from the IMU positioned in close proximity to the camera during the expedition. A few clearly distinguishable floes were manually segmented after orthorectification by outlining their contours using photo editing software for analysis (see Fig. 3).
@@ 3.3 Image enhancement
For night photographs, enhancement was limited to areas illuminated by the vessel’s spotlight. These regions are identified using haze removal techniques (He et al., 2011), which estimate the illumination map by inverting the low-light image, applying dehazing, and reinverting the result. This highlights the spotlight-illuminated areas for further analysis. As with daytime images, segmentation performance in these illuminated areas is influenced by sea ice complexity.
@@ 4 Image segmentation
4.1 Active contours model (GVF)
@@
The adjacent objects are then separated by smoothing their edges through a morphological erosion process (Adams, 1993). The newly formed objects are initialized using ellipses, which approximate the shape of ice floes (Alberello et al., 2019; Passerotti et al., 2022). The initial contour is then chosen such that the second moment of the covariance matrix (a $\times2$ matrix that captures the object’s spatial spread and orientation) aligns with that of the object (Mulchrone and Choudhury, 2004) (see Fig. 4d).
@@
The GVF snake algorithm can produce anomalies and inaccuracies, known as “hallucinations”, where the segmentation detects features that do not correspond to actual floes. These occur when the algorithm misinterprets patterns or textures in the image—such as shadows, smooth gradients, or areas with low contrast—as floe boundaries. To identify and remove such spurious features, we evaluate the shape of each segmented object using coefficients of circularity and eccentricity, which quantify how closely contours resemble expected floe shapes. By applying deliberately high-valued thresholds, spurious features in the form of nearly perfect circles or line segments are removed (circularity >1 and eccentricity $\geq$0.9). These thresholds target segmentation noise across all images, such as isolated pixels, without affecting segmented floe masks. Additionally, objects with open boundaries at the image’s edges are excluded from the segmentation.
@@
From the final binary mask, each ice floe is identified as a connected region of pixels using eight-neighborhood connectivity, meaning that two pixels are considered connected if they share an edge or a corner. Once the floes are identified, geometric properties such as area, perimeter, and diameter are derived. The area is determined by counting the total number of pixels in each region. The perimeter is obtained by tracing along the region boundary to measure its continuous outline (Biswas and Hazra, 2018). The diameter is calculated by equating the floe to a circle of the same area, thus representing the diameter of that hypothetical circle. The ice concentration is computed by summing the areas of all floes and dividing by the total image size excluding any masked-out, empty regions introduced by orthorectification. We note that ice concentration is used here only as a relative metric for comparison with the manual benchmark, which similarly excludes floes partially located at image edges. For absolute concentration estimates, partial floes at the image edges should be included.
@@ 4.2 Segment Anything Model
Figure 5: SAM segmentation modes applied to enhanced sea ice imagery. (a–c) Automatic segmentation mode (SAM-A): (a) grid-based fixed points automatically selected by SAM; (b) consolidation of masks returned by the model after filtering; (c) final unified binary mask of detected floes. (d–f) Prompt-based segmentation mode (SAM-P): (d) seed points (red dots) marking floe locations; (e) consolidation of masks returned by the model after filtering; (f) final unified binary mask of detected floes.
@@ 4.2.1 Automatic mask generator (SAM-A)
SAM cannot differentiate between the various segmented elements and, hence, segmentation includes both ice floes and open water fractions. The latter is excluded by applying the convexity coefficient (Sklansky, 1970) as a filtering criterion. The complexity coefficient quantifies how closely an object’s boundary resembles that of a convex shape. Ice floes, with regular boundaries that can often be approximated by ellipses or circles (Alberello et al., 2019), typically exhibit convexity values close to 1. In contrast, open water regions, with more irregular boundaries, have lower convexity values. A cut-off value of 0.8 is set, such that masks with a convexity less than or equal to 0.8 are excluded (cf. Wang et al., 2024). This threshold represents the minimum value that reliably removes water regions while preserving genuine floe masks across all images. The remaining masks (floes) are analyzed to identify those with anomalous shapes or open boundaries at the edges of the image using eccentricity and circularity metrics (see §4.1). Masks exhibiting these irregularities are also eliminated because they are either inconsistent with typical floe geometry or incompletely segmented due to edge effects.
@@ 4.2.2 Prompt-based (SAM-P)
In prompt-based mode, segmentation is guided by user-defined seeds, which are-derived from the distance transform matrix introduced in §4.1 (Fig. 4b). These seeds pre-identify the floes, and, for each one, SAM-P generates a corresponding mask (Fig. 5d). In regions with poorly defined boundaries, the assignment of seeds to floes can be ambiguous, and the resulting mask does not always isolate the intended floe accurately. In such cases, the mask includes unwanted areas, such as discontinuous water regions or parts of adjacent floes (see the first two layers in Fig. 5e). Shape-based filters for circularity, eccentricity, and convexity, with user-defined thresholds (see §4.1 and §4.2.1), are applied to remove clusters of feature pixels that do not match the typical shape of ice floes. A union operation is also used to merge clusters representing parts of adjacent floes with the correct floe. This operation combined all the masks in the output list into a single unified mask, automatically consolidating the segmented portions of floes into their appropriate masks (Fig. 5f). From this final binary mask, the properties of each individual ice floe and sea ice concentration are extracted.